Question 1:
Part a: Evaluate a 4 - year project costing $25,000 and returning $8,000 annually using the payback period technique and a 3-year cutoff.
Answer: 3.125 Years

Part b: Evaluate the project in Problem #1 using the profitability index method and a 10% required return.

Part c: Evaluate the project in part "a" using the IRR method and a 10% required rate

Please show work and formula used
Answer #1

fc =IRR(13:17) I J K 1 Project B year Cash flows pv@10% Present value Cumulative Cash flows 0 $ 25,000.00) 1.0000 $ (25,000.0

18 fx =IRR(13:17) H pv@10% 1 Project B 2 year 3 0 4 1 52 6 3 74 IRR Cash flows -25000 8000 8000 8000 8000 =IRR(13:17) =3/1.1